OpenStudy (kaylardz0405):

If x can be any number, how many solutions are there for the equation? y = 3x – 1 A. There are many solutions because there are many values for the variables that make the equation true. B. There are two solutions because there are two variables in the equation. C. There is only one solution because all equations have one solution. D. There is no solution because there are no values for the variables that make the equation true.

OpenStudy (kaylardz0405):

@AmberAlexis

OpenStudy (kaylardz0405):

@jagr2713 @jim_thompson5910 @Nnesha

OpenStudy (anonymous):

I say A

OpenStudy (kaylardz0405):

why

OpenStudy (anonymous):

C. False, if x= 2, y = 9 and if x = 1, y = 8. D. False, if x= 2, y = 9 A. True, there is an infinity of values because of the definition domain of the equation ; y \in \mathbb {R}, so D_f = ]- \infty ; + \infty [ B. False, it's just a two variables equation (system)

OpenStudy (anonymous):

Sorry out of order

OpenStudy (kaylardz0405):

it all right

OpenStudy (anonymous):

Hope it helps

OpenStudy (kaylardz0405):

do yall agree